# From No-Regret to Strategically Robust Learning in Repeated Auctions

### v1

@@ Appendix B A high-swap-regret instance for agile OGD
We note that there is a growing body of work on stronger notions of regret that agile OGD minimizes (e.g., Ahunbay (2024), Ahunbay and Bichler (2025), Cai et al. (2025)), and it is known that agile OGD does not minimize swap regret (see e.g., Ahunbay and Bichler (2025)). However, we were unable to find in the literature an explicitly written example showing that agile OGD with a fixed step size can incur $\Omega(T)$ swap regret. For completeness, we formalize one such example here and hope that it may be useful for future research. We first define swap regret and present agile OGD (Algorithm 2)––more precisely, agile online gradient ascent, since we work with reward vectors rather than loss vectors––within the standard online learning setting (Definition 4.1).
@@
From Table 1, only Phase 2 contributes to this sum, because $r_{2}^{(t)}-r_{1}^{(t)}=1$ in Phase 2 and equals 0 in Phases 1 and 3. Hence, we have
@@
As long as the third component remains nonnegative, i.e.,
@@
Starting from $\bm{\pi}^{(t)}=(0,0,1)$ at the beginning of Phase 1 of each batch (we will explain later why every batch begins with this state, except the very first batch), we have $x_{t}=0$ and hence
@@
Since $\alpha=\frac{100}{\eta}>\frac{3}{2\eta}$, Phase 1 contains more than enough rounds to reach $(\tfrac{1}{2},\tfrac{1}{2},0)$. Observe that this point is a fixed point of the projection update under $\mathbf{r}^{(t)}=(1,1,0)$. Hence the state remains $(\tfrac{1}{2},\tfrac{1}{2},0)$ for the remainder of Phase 1.
For the very first batch, the same recurrence applied to the initial distribution $(\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3})$ drives it to $(\tfrac{1}{2},\tfrac{1}{2},0)$ within $O\big(\frac{1}{\eta}\big)$ rounds, well before Phase 1 ends.
Phase 2: $\mathbf{r}^{(t)}=(0,1,0)$.
At the start of Phase 2 in any batch, we are at $\bm{\pi}^{(t)}=(\tfrac{1}{2},\tfrac{1}{2},0)$. During this phase, the state stays on the edge $\{(\pi_{1},\pi_{2},0):\pi_{1}+\pi_{2}=1\}$. Write $\pi_{1}^{(t)}=x_{t}$, so $\pi_{2}^{(t)}=1-x_{t}$. We update
@@
$2(a-x_{t})+2(a-(x_{t}-\eta))=0\quad\Rightarrow\quad a=x_{t}-\frac{\eta}{2}.$
@@
and after $\frac{1}{\eta}<\alpha$ steps, $x_{t+\frac{1}{\eta}}$ hits 0, and the state becomes $(0,1,0)$ and stays there for the remainder of Phase 2.
@@
Phase 3: $\mathbf{r}^{(t)}=(0,0,1)$.
At the start of Phase 3, we are at $\bm{\pi}^{(t)}=(0,1,0)$. During this phase, the state stays on the edge $\{(0,\pi_{2},\pi_{3}):\pi_{2}+\pi_{3}=1\}$. Write $\pi_{2}^{(t)}=x_{t}$, so $\pi_{3}^{(t)}=1-x_{t}$.
We update
@@
Starting from $x_{t}=1$, after $\frac{2}{\eta}<\alpha$ steps we reach $(0,0,1)$. Hence, each Phase 3 ends at $(0,0,1)$. Therefore, every new batch starts from $(0,0,1)$, and the above analysis of Phases 1 and 2 repeats identically in each batch.
Total swap regret.

### v2

@@ Appendix B A high-swap-regret instance for agile OGD
We note that there is a growing body of work on stronger notions of regret that agile OGD minimizes (e.g., Ahunbay (2024), Ahunbay and Bichler (2025), Cai et al. (2025)), and it is known that agile OGD does not minimize swap regret (see e.g., Ahunbay and Bichler (2025)). However, we were unable to find in the literature an explicitly written example showing that agile OGD with a fixed step size can incur $\Omega(T)$ swap regret. For completeness, we formalize one such example here and hope that it may be useful for future research. We first define swap regret and present agile OGD (Algorithm 2)—more precisely, agile online gradient ascent, since we work with reward vectors rather than loss vectors—within the standard online learning setting (Definition 4.1).
@@
From Table 1, only Phase 2 contributes to this sum, because $r_{2}^{(t)}-r_{1}^{(t)}$ equals 1 in Phase 2 and equals 0 in Phases 1 and 3. Hence, we have
@@
As long as the third component remains non-negative, i.e.,
@@
Starting from $\bm{\pi}^{(t)}=(0,0,1)$ at the beginning of Phase 1 of each batch (in the analysis of Phase 3, we will explain why every batch begins with this state, except the very first batch), we have $x_{t}=0$ and hence
@@
Since $\alpha=\frac{100}{\eta}>\frac{3}{2\eta}$, Phase 1 contains more than enough rounds to reach $(\tfrac{1}{2},\tfrac{1}{2},0)$. Observe that this point is a fixed point under the update rule of Algorithm 2 when $\mathbf{r}^{(t)}=(1,1,0)$. Hence the state remains $(\tfrac{1}{2},\tfrac{1}{2},0)$ for the remainder of Phase 1.
Analogously, for the very first batch, the same recurrence applied to the initial distribution $(\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3})$ reaches $(\tfrac{1}{2},\tfrac{1}{2},0)$ within $\alpha=\frac{100}{\eta}$ rounds.
Phase 2: $\mathbf{r}^{(t)}=(0,1,0)$
At the start of Phase 2 in any batch, the iterate is $\bm{\pi}^{(t)}=(\tfrac{1}{2},\tfrac{1}{2},0)$. During this phase, the state stays on the edge $\{(\pi_{1},\pi_{2},0):\pi_{1}+\pi_{2}=1\}$. Suppose that $\pi_{1}^{(t)}=x_{t}$, and hence, $\pi_{2}^{(t)}=1-x_{t}$. We update
@@
$2(a-x_{t})+2(a-(x_{t}-\eta))=0\quad\Longrightarrow\quad a=x_{t}-\frac{\eta}{2}.$
@@
and after $\frac{1}{\eta}<\alpha$ steps, $x_{t+\frac{1}{\eta}}$ decreases to 0, and the state becomes $(0,1,0)$ and stays there for the remainder of Phase 2.
@@
Phase 3: $\mathbf{r}^{(t)}=(0,0,1)$
At the start of Phase 3, the iterate is $\bm{\pi}^{(t)}=(0,1,0)$. During this phase, the state stays on the edge $\{(0,\pi_{2},\pi_{3}):\pi_{2}+\pi_{3}=1\}$. Suppose that $\pi_{2}^{(t)}=x_{t}$, and hence, $\pi_{3}^{(t)}=1-x_{t}$. We update
@@
Starting from $x_{t}=1$, after $\frac{2}{\eta}<\alpha$ steps, the iterate reaches $(0,0,1)$. Hence, each Phase 3 ends at $(0,0,1)$. Therefore, every new batch starts from $(0,0,1)$, and the above analysis of Phases 1 and 2 repeats identically in each batch.
Total swap regret